OpenStudy (anonymous):

x-100+x=180 (how do you solve for x?)

OpenStudy (lgbasallote):

x - 100 + x = 180 combine the x's 2x - 100 = 180 add 100 to both sides 2x = 180 + 100 combine like terms 2x = 280 solve for x does that help? or do you need more help?

OpenStudy (anonymous):

Line segments and intersect at point E. What is the measure of in degrees?